-2(2y-7)=2y-12+6y I am lost please help
Answer by edjones(8007)

-2(2y-7)=2y-12+6y
-4y+14=2y-12+6y
-4y+14=8y-12
-12y=-26
y=26/12
y=13/6
.
Got it?
.
Ed